Assume that you want to construct a 95% confidence interval estimate of a population
mean.Find the estimate of the sample size needed to obtain the specified margin of error
for the 95% confidence interval. The sample standard deviation is given below:
. (Round to the
Margin of error is $7, standard deviation is $24.The required sample size is_
nearest whole number).
1
Expert's answer
2021-07-26T05:57:39-0400

The following expression to compute the confidence interval for the mean is used:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}},\bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}}),"

where the critical value correspond to critical values associated to the Normal distribution. The critical values for the given "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96,"


"margin\ of\ error=E=z_c\times\dfrac{\sigma}{\sqrt{n}}"


Then


"n=(\dfrac{z_c\sigma}{E})^2"

Given "E=7, \sigma=24"


"n=(\dfrac{1.96(24)}{7})^2=45"

The required sample size is 45.